OpenStudy (anonymous):

Graph the equation by plotting points Y=1/4x-5 Complete the ordered pairs (4,_), (0,_)

OpenStudy (anonymous):

(4,1/11), (0,-1/5)

OpenStudy (anonymous):

u basically plug in 4 and 0 for wherever u have X then it gives u Y

OpenStudy (phi):

Is your equation \( y= \frac{1}{4x-5} \) or \( y=\frac{1}{4}x-5 \) or something else?